OpenStudy (anonymous):

Seven more than 2 times a number is 23.

OpenStudy (dangerousjesse):

7+2n=23

OpenStudy (dangerousjesse):

23-7=16 16\(\div\)2=8 n=8

OpenStudy (anonymous):

The difference of 2 times a number and 7 is 23.

OpenStudy (cj49):

let the number be "n".seven more means 7 plus and two times a number means 2x n....now can u slove it @THEUSER822

OpenStudy (anonymous):

7n + 2 = 23

OpenStudy (cj49):

nopes two times a number means 2x n = 2n therefore 7+2n=23

OpenStudy (dangerousjesse):

2n-7=23 23+7=30 30\(\div\)2=15 for the second one.

OpenStudy (dangerousjesse):

So n=15

OpenStudy (anonymous):

The sum of 7 times a number and 2 is 23.

OpenStudy (dangerousjesse):

What do you think? :)

OpenStudy (anonymous):

@DangerousJesse @ninamae05 @matimaticas

OpenStudy (anonymous):

i think 2n + 7 = 23

OpenStudy (dangerousjesse):

So you're saying "7 times a number"=2n?

OpenStudy (anonymous):

what about this one it my second one 7n + 2 = 23

OpenStudy (dangerousjesse):

Yes, that's correct. :)

OpenStudy (anonymous):

1 more Seven less than 2 times a number is 23.

OpenStudy (anonymous):

Seven less than 2 times a number is 23

OpenStudy (dangerousjesse):

x less than y= y-x What do you think?

OpenStudy (anonymous):

2n-7=23

OpenStudy (dangerousjesse):

Yes, that's right.

OpenStudy (anonymous):

thx